# PolyFly: Polytopic Optimal Planning for Collision-Free Cable-Suspended Aerial Payload Transportation

### v1

@@ I Introduction
In time critical scenarios, such as rescue missions, it is crucial for aerial vehicles to optimize their flight path to reduce the mission time, especially in challenging environments like dense forests or earthquake-damaged and structurally unsafe buildings. This requires the robot planners to avoid obstacles while maneuvering through narrow gaps and around tight corners, which need non-conservative physical approximations of the system components. Furthermore, unlike standalone quadrotor methods, aerial transportation planners must take into account the increased complexity of the configuration space introduced by the swinging payload.
Existing methods often simplify these systems into spheres [6] or large prisms [7, 8] to address computational demands, trading off solution quality for computational speed. In contrast, our approach directly addresses these limitations by representing each component - quadrotor, payload, and cable — as distinct polytopes. We further enhance our planner by incorporating the quadrotor’s orientation into the polytopic representation. This orientation-aware modeling enables more precise collision prediction during aggressive maneuvers in confined spaces such as the environment shown in Fig. 4. This representation provides planners with detailed geometric information, enhancing their capabilities to navigate maze-like and highly constrained environments.
Current aerial transportation planning methodologies can broadly be categorized into sampling-based and optimization-based approaches. Within optimization methods, trajectories are either parameterized through polynomial representations [6, 9] or discretized and solved via Model Predictive Control (MPC) frameworks [10, 11, 12, 7]. Our work builds upon MPC-based global planning techniques and specifically addresses the limitations posed by existing approaches that simplify system geometry. Methods that employ Euclidean Signed Distance Fields (ESDFs) result in discontinuous optimization constraints [13] while cylindrical obstacle representations limit the complexity of environments that can be modeled. We tackle these issues by formulating obstacle avoidance explicitly through polytopic models and represent each robot component and every obstacle as distinct polytopes. As we show in this work, this representation allows us to both incorporate smooth differentiable constraints into our global planner, and generate aggressive trajectories in tightly constrained environments.
@@
  • A novel non-conservative representation for aerial transportation that models the quadrotor, cable, and payload as separate polytopes. To enhance physical accuracy, we incorporate the attitude of both the robot and the obstacles into the environment representation, constructing orientation-aware polytopes that more accurately capture the system’s spatial footprint.
@@
  • Experimental validation on hardware by performing real-world experiments using a quadrotor carrying a slung payload. We illustrate the advantages of our representation and planning method by tracking trajectories in tightly constrained environments. We also demonstrate that our approach generates faster trajectories than the existing state-of-the-art method in all eight test environments.
We plan to release our implementation as an open-source standalone package with minimal dependencies. Unlike other available frameworks that are integrated into extensive simulation environments, our package facilitates straightforward benchmarking and comparison, lowering the barrier to entry and accelerating innovation in planning and control for aerial transportation.
@@ II Related Works
In this section, we survey the main design choices in (i) how the environment and robots are spatially represented and (ii) how the states and inputs are mathematically modeled. We emphasize advancements that leverage polytopic geometry and highlight the benefits of PolyFly over prior works.
@@
Environment Representations. Obstacle-aware planners typically encode the workspace either implicitly, through Exponential Signed Distance Fields (ESDFs), or explicitly, by representing obstacles with geometric shapes. Recent aerial transportation planners [6, 9], voxelize the environment and store an ESDF, turning collision checks into fast table look-ups. Although highly expressive, ESDFs can introduce non-smooth, discontinuous gradients, especially near object boundaries, affecting convergence in non-linear programs. As discussed in Section V-Q1, we believe this is one of the reasons such methods fail to converge and are outperformed by PolyFly.
@@
Other works express the environment via geometric objects. [11, 12, 10] employ cylindrical and spherical objects, while [7, 8, 14] rely on polytopes. While ellipses and cylinders limit the range of scenes that can be modeled, polytopes significantly increase the complexity of environments that can be captured. In fact, works that use ESDFs also primarily consider environments composed of multiple cuboids [9]. Furthermore, in the context of trajectory planning, adopting polytopes to represent the environment yields distance constraints with good optimization properties. Duality theory can be used to convert non-linear polytope–to-polytope distance inequalities into smooth, differentiable constraints. [15, 16] used polytopic collision constraints for short-horizon reactive MPC frameworks but are unable to optimize over global paths due the short MPC horizon. [7] used this approach for global planning but did not consider complex environments and cannot navigate environments through narrow gaps, eg. Fig. LABEL:fig:cover_page, due to their conservative representation that models the robot as a single cuboid.
@@
Figure 1: System diagram showing the four different polytopes used in PolyFly: Quadrotor $N_{Q}$, Payload $N_{L}$, Cable $N_{C}$, and Obstacles $M_{i}$. The quadrotor polytope enables accurate distance computations ${h}_{N_{Q},M_{i}}$, whereas most methods use an ellipse around the quadrotor resulting in conservative and inaccurate distance computations $\tilde{h}_{N_{Q},M_{i}}$.
@@
Spatial Robot Representation. There are two primary design decisions with respect to modeling the robot. First, how should the physical geometry of the robot be modeled? Existing methods typically use multiple spheres or ellipses to represent the quadrotor and its payload [6, 9, 11, 12, 10] or a single prism or polytope [7, 8]. These representations over-approximate at least one component (quadrotor, payload or cable), therefore limiting their applicability in highly constrained or cluttered environments. In contrast, our method addresses these limitations by leveraging orientation-aware polytopes for each robot component - quadrotor, cable, and payload. Combined with polytopic representation of the environments, we enable non-conservative distance evaluation, as shown in Fig. 1, facilitating aerial transportation in maze-like environments and narrow passages that are slightly larger than the width of the cable (see Fig. 4).
@@
State and Input Modeling. The second design decision is how states and inputs are mathematically represented. A popular strategy is to express the states and inputs as polynomials to facilitate trajectory optimization [6, 9]. This is often paired with MINCO [17] to reformulate the problem into an unconstrained optimization to boost computational efficiency. While this approach can be used to achieve real-time capabilities, our experiments in Section V-Q1 indicate that it may compromise both solution quality and convergence. In contrast, similar to [11, 7, 12, 10], PolyFly operates directly in the native state space. Unlike polynomial parameterizations, the native state space approach places no restrictions on the shape of the state or control trajectories, making them inherently more expressive. Our experiments confirm that this added flexibility enables PolyFly to produce faster trajectories and tackle more complex environments than the state-of-the-art polynomial-based planner.
@@
The methods presented in [6, 9], which serve as our baseline, employ an ESDF to represent the environment, approximate the quadrotor and payload with spheres, and parameterize the states and inputs via polynomials. [9] reports an ablation study comparing the computational times and success rates of their method with IPOPT-MP, a state-based global planning approach. However, they omit the comparison of the optimized trajectory times of their method with IPOPT-MP, and do not address navigation close to obstacles in maze-like environments. We demonstrate that by using the native state-based modeling and a non-conservative spatial workspace representation via polytopes, PolyFly outperforms the baseline in all test environments.
@@ III System Dynamics
$m_{L},m_{Q}$ mass of payload and robot
$\mathbf{x}_{L}{},\mathbf{x}_{Q}{}$ position of payload and robot in $\mathcal{I}$
$\dot{\mathbf{x}}_{L},\ddot{\mathbf{x}}_{L}$ linear velocity, acceleration of payload in $\mathcal{I}$
$\dddot{\mathbf{x}}_{L}$ jerk of payload in $\mathcal{I}$
$\dot{\mathbf{x}}_{Q},\ddot{\mathbf{x}}_{Q}$ linear velocity, acceleration of robot in $\mathcal{I}$
@@
$\mathbf{\Omega}{}\in\mathbb{R}^{3}$ angular velocity of robot in $\mathcal{B}{}$
@@
$f_{i}\in\mathbb{R}$ $i^{\text{th}}$ motor thrust
$J\in\mathbb{R}^{3\times 3}$ moment of inertia of robot
@@
$l,g\in\mathbb{R}$ cable length, gravity constant
$\mathbf{q}$ quaternion representation of $\bm{\mathit{R}}_{Q}{}$
@@
$\frac{d\mathbf{x}_{L}}{dt}=\dot{\mathbf{x}}_{L},~\frac{d\mathbf{x}_{Q}{}}{dt}$ $=\dot{\mathbf{x}}_{Q}{},~\dot{\mathbf{q}}=\frac{1}{2}\hat{\mathbf{\Omega}{}}\mathbf{q},$ (1)
@@
$\begin{bmatrix}f\\[2.0pt] \mathbf{M}\end{bmatrix}$ $=\mathbf{D}\begin{bmatrix}f_{1}\\ f_{2}\\ f_{3}\\ f_{4}\end{bmatrix}$ (5)
where $\mathbf{g}=g\mathbf{e}_{3}$, $g=9.81$ $m/s^{2}$, $\mathbf{e}_{3}=\left[0~0~1\right]^{\top}$, and $\hat{\mathbf{\Omega}{}}$ is the skew-symmetric matrix of the quadrotor angular velocity $\mathbf{\Omega}{}$. Table I defines the remaining variables. The matrix $\mathbf{D}$ provides a linear mapping between collective thrust-moments and individual motor thrusts. It is defined in terms of the robot’s physical properties and propeller’s aerodynamics characteristics. Readers can refer to [18] for additional information.
By leveraging the differential flatness property of the system [3], we can model the state dynamics described in eqs. (1)–(5) in terms of the flat outputs $\{\mathbf{x}_{L}{},\psi\}$, where $\psi$ is the quadrotor’s yaw angle. In our method, we make the non-restrictive assumption and set $\psi=0$. We choose to express the resulting dynamics as
$\mathbf{x}$ $=\begin{bmatrix}\mathbf{x}_{L}{}^{\top},\dot{\mathbf{x}}_{L}{}^{\top},\ddot{\mathbf{x}}_{L}{}^{\top}\end{bmatrix}^{\top},\mathbf{u}=\begin{bmatrix}\dddot{\mathbf{x}}_{L}^{\top}\end{bmatrix}^{\top}$. (6)
The induced system dynamics are therefore modeled as a third-order integrator. Given $\mathbf{x}$, $\mathbf{u}$ and the robot’s physical parameters, we can use the non-linear differential flatness mapping [3] to calculate the robot position $\mathbf{x}_{Q}{}$, robot velocity $\dot{\mathbf{x}}_{Q}{}$, robot acceleration $\ddot{\mathbf{x}}_{Q}{}$ and robot orientation $\bm{\mathit{R}}_{Q}{}$. This allows our trajectory generator to plan in payload space and add constraints to the quadrotor’s states. Since our planner operates in a discrete space, we convert the continuous time representation above to its discrete form. We obtain the Runge-Kutta approximation
$\mathbf{x}(k+1)=F(\mathbf{x}(k),\mathbf{u}(k),\Delta t_{k}),$ (7)
where $F$ represents the discretized dynamics function and $\Delta t_{k}$ the time step duration at stage $k$.
@@ IV-A Preliminaries
$\bm{\tau}$ $={}\begin{bmatrix}\bm{\tau}_{0},\bm{\tau}_{1},\dots,\bm{\tau}_{N}\end{bmatrix}$, (8)
$\bm{\tau}_{k}$ $={}\begin{bmatrix}\mathbf{x}_{L}(k),\,\dot{\mathbf{x}}_{L}(k),\,\ddot{\mathbf{x}}_{L}(k)\end{bmatrix}^{\top},~k\in(0,N).$
The robot starts at state $\mathbf{x}_{o}$ and must navigate to its goal state $\mathbf{x}_{N}$. The environment is described by the set $\mathbf{E}$, which contains $N_{O}$ polytopic obstacles $M_{i}$ for $i\in(0,N_{O})$. The aerial transportation system is modeled using three separate polytopes, $N_{j}$ for $j\in\{Q,C,L\}$ that represent the quadrotor, cable, and load polytopes respectively, as shown in Fig. 1. The rest of this section discusses our choice of distance metric between the robot and obstacles, and a detailed overview of our optimization approach.
@@ IV-B Polytopic Approximation
Let us consider two polytopes $M_{i}$ and $N_{j}$, where $M_{i}$ and $N_{j}$ represent the obstacle and robot polytopes respectively, as defined in Section IV-A. Let $\mathbf{y}_{M_{i}}$ be any point within the polytope $M_{i}$ and $\mathbf{y}_{N_{j}}$ be any point within the polytope $N_{j}$. $A_{M_{i}},B_{M_{i}},A_{N_{j}}$ and $B_{N_{j}}$ are constants that define their geometric shape. These polytopes can be represented by the standard inequalities
@@
where $\bm{\mathit{R}}_{N_{j}}$ is $N_{j}$}’s rotation matrix, and $\mathbf{O}_{N_{j}}$ is it’s origin. Substituting this into the polytopic form expressed by eq. (9), we can represent polytope $M_{i}$ as
@@
$(\mathbf{A}_{M_{i}}\bm{\mathit{R}}_{N_{j}}(\mathbf{x}))\mathbf{y}_{M_{i}}^{N_{j}}$ $\leq\mathbf{B_{M_{i}}}-\mathbf{A}_{M_{i}}\mathbf{O}_{N_{j}}(\mathbf{x}),$ (12)
$=\mathbf{A}_{M_{i}}^{N_{j}}(\mathbf{x})\mathbf{y}_{M_{i}}^{N_{j}}$ $\leq\mathbf{B}_{M_{i}}^{N_{j}}(\mathbf{x}),$ (13)
where $\mathbf{A}_{M_{i}}^{N_{j}}(\mathbf{x})=\mathbf{A}_{M_{i}}\bm{\mathit{R}}_{N_{j}}(\mathbf{x})$ and $\mathbf{B}_{M_{i}}^{N_{j}}(\mathbf{x})=\mathbf{B_{M_{i}}}-\mathbf{A}_{M_{i}}\mathbf{O}_{N_{j}}(\mathbf{x})$.

### v2

@@ I Introduction
In time critical scenarios, such as rescue missions, it is crucial for aerial vehicles to optimize their flight path to reduce the mission time, especially in challenging environments like dense forests or earthquake-damaged and structurally unsafe buildings. This requires the robot planners to avoid obstacles while maneuvering through narrow gaps and around tight corners, which need non-conservative physical approximations of the system components.
Existing methods often simplify these systems into spheres [6] or large prisms [24, 16] to address computational demands, trading off solution quality for computational speed. In contrast, our approach directly addresses these limitations by representing each component - quadrotor, payload, and cable - as distinct polytopes. We further enhance our planner by incorporating the quadrotor’s orientation into the polytopic representation. This orientation-aware modeling enables more precise collision prediction during aggressive maneuvers in confined spaces such as the environment shown in Fig. 4. This representation provides planners with detailed geometric information, enhancing their capabilities to navigate maze-like and highly constrained environments.
Current aerial transportation planning methodologies can broadly be categorized into sampling-based and optimization-based approaches. Within optimization methods, trajectories are either parameterized as polynomials [6, 20] or discretized and solved via Model Predictive Control (MPC) frameworks [14, 4, 13, 24]. Our work builds upon MPC-based global planning techniques and specifically addresses the limitations posed by existing approaches that simplify system geometry. Methods that employ Euclidean Signed Distance Fields (ESDFs) result in discontinuous optimization constraints [11] while cylindrical obstacle representations limit the complexity of environments that can be modeled. We tackle these issues by formulating obstacle avoidance explicitly through polytopic models and represent each robot component and every obstacle as distinct polytopes. As we show in this work, this representation allows us to both incorporate smooth differentiable constraints into our global planner, and generate aggressive trajectories in tightly constrained environments.
@@
  • A novel non-conservative representation for aerial transportation that models the quadrotor, cable, and payload as separate components. To enhance physical accuracy, we incorporate the attitude of both the robot and the obstacles into the environment representation, constructing orientation-aware polytopes that more accurately capture the system’s spatial footprint.
@@
  • Experimental validation on hardware by performing real-world experiments using a quadrotor carrying a slung payload. We illustrate the advantages of our representation and planning method by tracking trajectories in tightly constrained environments. We also demonstrate that our approach generates faster trajectories than the existing state-of-the-art method in all ten test environments.
We release our implementation as an open-source, standalone package with minimal dependencies.
@@ II Related Works
In this section, we survey the main design choices in (i) how the environment and robots are spatially represented and (ii) how the states and inputs are mathematically modeled.
@@
Environment Representations. Obstacle-aware planners typically encode the workspace either implicitly, through Euclidean Signed Distance Fields (ESDFs), or explicitly, by representing obstacles with geometric shapes. Recent aerial transportation planners [6, 20], voxelize the environment and store an ESDF, turning collision checks into fast table look-ups, but introduce non-smooth, discontinuous gradients when used in optimization-based solvers.
@@
Other works express the environment via geometric objects. [4, 13, 14] employ cylindrical and spherical objects, while [24, 16, 23] rely on polytopes. While ellipsoids and cylinders limit the range of scenes that can be modeled, polytopes significantly increase the complexity of environments that can be captured. In fact, works that use ESDFs also primarily consider environments composed of multiple cuboids [20]. Furthermore, in the context of trajectory planning, adopting polytopes to represent the environment yields distance constraints with good optimization properties. Duality theory can be used to convert non-linear polytope–to-polytope distance inequalities into smooth, differentiable constraints. [8, 17] used polytopic collision constraints for short-horizon reactive MPC frameworks but are unable to optimize over global paths due to the short MPC horizon. [24] used this approach for global planning but did not consider complex environments and cannot navigate environments through narrow gaps, eg. Fig. LABEL:fig:cover_page, due to their conservative representation that models the robot as a single cuboid.
@@
Figure 1: Different polytopes used in PolyFly: Quadrotor $N_{Q}$, Payload $N_{L}$, Cable $N_{C}$, and Obstacles $M_{i}$. The quadrotor polytope enables accurate distance computations ${h}_{N_{Q},M_{i}}$, whereas most methods use a sphere around the quadrotor resulting in conservative distance computations $\tilde{h}_{N_{Q},M_{i}}$.
@@
Spatial Robot Representation. There are two primary design decisions with respect to modeling the robot. First, how should the physical geometry of the robot be modeled? Existing methods typically use multiple spheres or ellipses to represent the quadrotor and its payload [6, 20, 4, 13, 14] or a single prism or polytope [24, 16]. These representations over-approximate at least one component (quadrotor, payload, cable or environment obstacles), therefore limiting their applicability in highly constrained or cluttered environments. Spheres over-approximate the quadrotor, cable, and obstacles while ellipses over-approximate the quadrotors and obstacles used in this work. For example, ellipsoidal representations fail in the ceiling-wall environments in Fig. 6, while single polytope representations [24] fails in the ceiling-floor and narrow gap environments in Fig. 4 and 6. In contrast, our method addresses these limitations by leveraging orientation-aware polytopes for each robot component - quadrotor, cable, and payload. Combined with polytopic representation of the environments, we enable non-conservative distance evaluation, as shown in Fig. 1, facilitating aerial transportation in maze-like environments and narrow passages that are slightly larger than the width of the cable.
@@
State and Input Modeling. The second design decision is how states and inputs are mathematically represented. A popular strategy is to express the states and inputs as polynomials to facilitate trajectory optimization [6, 20]. This is often paired with MINCO [21] to reformulate the problem into an unconstrained optimization to boost computational efficiency. While this approach can be used to achieve real-time capabilities, our experiments in Section V-Q1 indicate that it may compromise both solution quality and convergence. In contrast, similar to [4, 24, 13, 14], PolyFly operates directly in the native state space. Unlike polynomial parameterizations, the native state space approach places no restrictions on the shape of the state or control trajectories, making them inherently more expressive.
@@
The method presented in [6, 20], serves as our primary baseline, employs an ESDF to represent the environment, approximates the quadrotor and payload with spheres, and parameterizes the states and inputs via polynomials. [20] reports an ablation study comparing the computational times and success rates of their method with IPOPT-MP, a state-based global planning approach. However, they omit the comparison of the optimized trajectory times of their method with IPOPT-MP, and do not address navigation close to obstacles in maze-like environments. We demonstrate that by using the native state-based modeling and a non-conservative spatial workspace representation via polytopes, PolyFly outperforms this method in all test environments.
@@ III System Dynamics
$m_{L},m_{Q}\in\mathbb{R}$ mass of payload and robot
$\mathbf{x}_{L}{},\mathbf{x}_{Q}{}\in\mathbb{R}^{3}$ position of payload and robot in $\mathcal{I}$
$\dot{\mathbf{x}}_{L},\ddot{\mathbf{x}}_{L},\dddot{\mathbf{x}}_{L}\in\mathbb{R}^{3}$ payload velocity, acceleration, jerk in $\mathcal{I}$
$\dot{\mathbf{x}}_{Q},\ddot{\mathbf{x}}_{Q}\in\mathbb{R}^{3}$ robot velocity, acceleration in $\mathcal{I}$
@@
$\mathbf{\Omega}{}\in\mathbb{R}^{3}$ robot angular velocity in $\mathcal{B}{}$
@@
$J\in\mathbb{R}^{3\times 3}$ moment of inertia of robot in $\mathcal{B}{}$
@@
$f_{i},l,g\in\mathbb{R}$ $i^{\text{th}}$ motor thrust, cable length, gravity
$N_{O}\in\mathbb{Z}$ total number of obstacles
$M_{i}$, $N_{j}$ $i^{th}$ obstacle and $j^{th}$ robot polytope,
@@
$\frac{d\mathbf{x}_{L}}{dt}=\dot{\mathbf{x}}_{L},~\frac{d\mathbf{x}_{Q}{}}{dt}$ $=\dot{\mathbf{x}}_{Q}{},~\dot{\mathbf{R}}_{Q}=\bm{\mathit{R}}_{Q}{}\hat{\mathbf{\Omega}}$, (1)
@@
$\begin{bmatrix}f,\mathbf{M}\end{bmatrix}^{\top}$ $=\mathbf{D}\begin{bmatrix}f_{1},f_{2},f_{3},f_{4}\end{bmatrix}^{\top}$. (5)
where $\mathbf{g}=g\mathbf{e}_{3}$, $g=9.81$ $m/s^{2}$, $\mathbf{e}_{3}=\left[0~0~1\right]^{\top}$, and $\hat{\mathbf{\Omega}{}}$ is the skew-symmetric matrix of the quadrotor angular velocity $\mathbf{\Omega}{}$. Table I defines the remaining variables. Qualitatively, eq. (2) describes the relationship between the payload acceleration and collective motor thrust, while eq. (3) shows that cable’s angular acceleration depends on the collective thrust and the quadrotor’s orientation. The matrix $\mathbf{D}$ provides a linear mapping between collective thrust-moments and individual motor thrusts. It is defined in terms of the robot’s physical properties and propellers aerodynamics characteristics. Readers can refer to [5] for additional information.
By leveraging the differential flatness property of the system [15], we can model the state dynamics described in eqs. (1)–(5) in terms of the flat outputs $\{\mathbf{x}_{L}{},\psi\}$, where $\psi$ is the quadrotor’s yaw angle. In our method, we set $\psi=0$ and assume negligible cable angular accelerations. This allows us to express the system through the state $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{L}{}^{\top},\dot{\mathbf{x}}_{L}{}^{\top},\ddot{\mathbf{x}}_{L}{}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{9}$ and input $\mathbf{u}=\begin{bmatrix}\dddot{\mathbf{x}}_{L}\end{bmatrix}\in\mathbb{R}^{3}$.
The induced system dynamics are therefore modeled as a third-order system. Given $\mathbf{x}$, $\mathbf{u}$, the robot’s physical parameters, and assuming that cable angular accelerations are zero at the states optimized by the planner, we can use the non-linear differential flatness mapping [15] to calculate the robot position $\mathbf{x}_{Q}{}$, robot velocity $\dot{\mathbf{x}}_{Q}{}$, and orientation $\bm{\mathit{R}}_{Q}{}$. This allows our trajectory generator to plan in payload space and add constraints to the quadrotor’s states. Since our planner operates in a discrete space, we convert the continuous time representation above to its discrete form. We obtain the Runge-Kutta approximation $\mathbf{x}(k+1)=F(\mathbf{x}(k),\mathbf{u}(k),\Delta t_{k})$, where $F$ is the discretized dynamics function and $\Delta t_{k}$ the time step duration at stage $k$.
@@ IV-A Preliminaries
$\bm{\tau}={}\begin{bmatrix}\bm{\tau}_{0},\dots,\bm{\tau}_{N}\end{bmatrix},\bm{\tau}_{k}={}\begin{bmatrix}\mathbf{x}_{L}(k),\,\dot{\mathbf{x}}_{L}(k),\,\ddot{\mathbf{x}}_{L}(k)\end{bmatrix}^{\top}$, (6)
where $k\in(0,N)$. The robot starts at state $\mathbf{x}_{O}$ and must navigate to its goal state $\mathbf{x}_{N}$. The environment is described by the set $\mathbf{E}$, which contains $N_{O}$ polytopic obstacles $M_{i}$ for $i\in(0,N_{O})$. The aerial system is modeled using three separate polytopes, $N_{j}$ for $j\in\{Q,C,L\}$ that represent the quadrotor, cable, and load polytopes respectively, as shown in Fig. 1.
@@ IV-B Polytopic Approximation
Let us consider two polytopes $\mathbf{M_{i}}$ and $\mathbf{N_{j}}$, as defined in Table I. Let $\mathbf{y}_{M_{i}}$ be any point within the polytope $\mathbf{M_{i}}$ and $\mathbf{y}_{N_{j}}$ be any point within the polytope $\mathbf{N_{j}}$. $\mathbf{A_{M_{i}}},\mathbf{B_{M_{i}}},\mathbf{A_{N_{j}}}$ and $\mathbf{B_{N_{j}}}$ are constants that define their geometric shape. These polytopes can be represented by
@@
where $\bm{\mathit{R}}_{N_{j}}$ is $N_{j}$}’s rotation matrix, and $\mathbf{O}_{N_{j}}$ is its origin. Substituting this into the polytopic form expressed by eq. (7), we can represent polytope $M_{i}$ as
@@
$\mathbf{A}_{M_{i}}^{N_{j}}(\mathbf{x})\mathbf{y}_{M_{i}}^{N_{j}}$ $\leq\mathbf{B}_{M_{i}}^{N_{j}}(\mathbf{x}),$ (10)
$\mathbf{A}_{M_{i}}^{N_{j}}(\mathbf{x})=\mathbf{A}_{M_{i}}\bm{\mathit{R}}_{N_{j}}(\mathbf{x})$ $,\mathbf{B}_{M_{i}}^{N_{j}}(\mathbf{x})=\mathbf{B_{M_{i}}}-\mathbf{A}_{M_{i}}\mathbf{O}_{N_{j}}(\mathbf{x}).$